# BlackboxBench: A Comprehensive Benchmark of Black-box Adversarial Attacks

### v1

@@ IV-B Result Overview
In query-based attacks, we evaluate 25 methods outlined in Sec. III-B using datasets and models detailed in Sec. IV-A across all four attack settings. $\ell_{2}$ and $\ell_{\infty}$ are implemented for most attacks, even though some attacks are explicitly designed for one certain norm. There are a total of 570 pairs of evaluations for decision-based black-box attacks and 680 pairs of evaluations for score-based black-box attacks. Each method’s performance in attacking a target model is assessed using three metrics: AQN, MQN, and ASR. A better attacker should report a higher ASR and a lower AQN and MQN simultaneously. In transfer-based attacks, we evaluate 30 methods mentioned in Sec. III-B on datasets and models listed in Sec. IV-A under all four attack settings, leading to 13,132 evaluations in total. The efficacy of adversarial perturbations crafted on a surrogate model by each method are measured by averaging ASR on all targeted models listed in Tab. III. A superior attack method is expected to yield a higher average ASR. For the context of transfer-based attacks, unless explicitly specified otherwise, "ASR" refers to the averaged ASR herein.
Due to space limitations, in Fig. 6 and Fig. 5, we exclusively offer a condensed summary of results specifically for the ImageNet dataset alongside one model architecture—ResNet-50, $i.e.$, we only exhibit results where ResNet50 acts as surrogate model in transfer-based attacks and as target model in query-based attacks. For comprehensive tables including additional models and CIFAR-10 dataset, please refer to BlackboxBench LeaderBoard. Based on the results in Fig. 6 and Fig. 5, our analysis in this section mainly focuses on three aspects: 1) the evolution of black-box adversarial attacks over the years and across the categories; 3) the comparison between targeted and untargeted attacks.
IV-B1 Evolution over time and across categories
The overview of decision-based attack performance is depicted in the top row of Fig. 5. At the early stage of this field, represented by Boundary attack [92] and OPT[97], a prominent characteristic is the low ASR coupled with a relatively high AQN. However, a turning point occurred post-2020, where emerging targeted attacks marked a shift with both improved ASR and decreased AQN, and untargeted attacks demonstrated an improvement in ASR without costing more queries. In the category of gradient estimation, methods rooted in the HSJA [100] framework (beginning with QEBA [4], followed by NonLinear-BA [5], and culminating in PSBA [6]) exhibited significant attack capabilities—a decline in AQN while aiming for peak ASR. Their continuous pursuit of methodological advancements has consequently positioned gradient estimation-based techniques as superior in $\ell_{\infty}$ attack performance when compared to random search-based approaches, as observed in Fig. 5 (c). Nevertheless, within $\ell_{2}$ norm attacks, random search-based methods, SFA [95] and Rays [96], have carved a niche for themselves by recording the best of ASR and the lowest of AQN.
The bottom row of Fig. 5 shows the result overview of scorer-based attacks in BlackboxBench. Between 2018 and 2019, methods like ZO-signSGD [13] and SimBA [7] had high AQNs but yielded lower ASRs. However, with time, newly proposed methods showed improvements: $\ell_{\infty}$ attacks showcased enhanced ASR alongside reduced AQN, meanwhile $\ell_{2}$ attacks exhibited improved ASR without an increase in query costs. Notably, Square attack [9], along with the recently introduced methods $\mathcal{CG}$-attack [18] and MCG [19], stood out by demonstrating the highest ASRs and the lowest AQNs. Analyzing between categories, despite random search-based methods and gradient estimation-based methods performing evenly, we notice the surprising achievement of combination-based methods, achieving a perfect ASR of 1 with fewer queries, highlighting the significant role played by the surrogate model’s assistance. Indeed, numerous contemporary score-based black-box attacks utilize the extra information about adversarial samples offered by the surrogate model. BlackboxBench will include more combination-based methods in the future.
The result overview for transfer-based attacks in BlackboxBench is shown in Fig. 6(a). As depicted in Fig. 6(b), there appears an obvious increasing trend of attackers’ performance during the past seven years, suggesting continuing efforts within the active research community. At the early stage, researchers favor the idea that the transferability of the adversarial examples is similar to the generalization ability of the trained models[26]. Accordingly, attempts are more made from data-transformation and gradient-optimization perspectives. Later, inspired by the fact that different DNN models share similar features, transfer-based attacks from a feature perspective have shown promise in relieving the overfitting to the white-box model and the low transferability to black-box model by performing attacks in the intermediate layers. Fig. 6(c) illustrates that, from both data and optimization perspectives, the methods display nearly equal competitiveness, while feature-based methods exhibit superiority in specific attack settings. Recently, there has been a growing focus on refining surrogate models to enhance transferability. This approach has garnered significant attention and has demonstrated exceptional performance, particularly in targeted scenarios, albeit at the cost of increased computational overhead due to additional training requirements.
IV-B2 Untargeted vs. targeted
@@ IV-B1 Evolution over time and across categories
Regarding attacks with different targets, we can see that targeted query-based attacks require more query costs than untargeted attacks while exhibiting a relatively lower ASR. Also achieving transferability in targeted attacks proves more challenging than in untargeted attacks. Despite our efforts to conduct evaluations with additional iterations and a logit loss following [67], along with a higher norm budget $\epsilon$ to boost targeted attack transferability, existing literature still shows its inability to fool models into predicting a targeted class, even though they perform well in untargeted attacks. Nevertheless, methods from the model perspective dramatically narrow the gap between targeted and untargeted tasks, illuminating a promising research direction to overcome the low transferability in targeted attack scenarios.
IV-C Attack Procedure Analysis
For a more qualitative analysis, we explore what is most salient to the target model when an image is under some attack. Using our analytical tools—FullGrad, we visualize the evolution of model attention maps on the perturbed images along the iteration of white-box attacking in transfer-based attacks, or the path traversed by query-based attackers. We are interested in both how successful attacks succeed and how failed attacks fail.
In query-based attacks, we choose an image from the category of squirrel, which is also selected to be successfully attacked by about $80\%$ of all methods. Following the similar presentation of FullGrad in transfer-based attacks, we show the attention maps through the attack process of unsuccessful attacks in Fig. 25 and successful attacks of both decision-based attacks and score-based attacks in Fig. 28 and Fig. 29. We can also observe that in these samples where attacks did not succeed, the target model continuously concentrated on the critical regions or nearby regions, which failed to mislead the target model and thus the target model can correctly identify them throughout the attack process. Conversely, during successful attacks, the adversarial images incrementally shift the model’s attention away from these critical regions (from the central part to another part). Eventually, the models are deceived into concentrating on insignificant areas, leading to the misclassification of the adversarial images as the category of a beaver.
IV-D Effect of Data
In this section, we investigate how attributes of data impact the black-box adversarial attack performance from two perspectives: 1) characteristic of hard samples, 2) input dimensionality.
IV-D1 Characteristic of Hard Samples

### v2

@@ V Conclusion
[…]
A Implemented Algorithms
Despite the continuous emergence of black-box adversarial attack methods, we adhere to two guidelines when selecting the methods to be incorporated into BlackboxBench. Firstly, from a chronological perspective, we select both traditional and the most recently proposed methods to illustrate the progression within this field. Secondly, from a categorical perspective, we opt for a diversity of methods that span the numerous sub-domains of black-box adversarial attacks, facilitating our evaluation of the developmental trajectory of different technical routes.
For the selected methods, we present an elaborate taxonomy of them in Fig. 3 in the main text. The first level of our taxonomy distinguishes three types of black-box adversarial attacks: decision-based query attacks, score-based query attacks, and transfer-based attacks. At the second level, we subdivide the existing literature within each type according to their technical routes. At the third level, we list the implemented attacks in our BlackboxBench. Given the division of the first level introduced in Sec. III-A, in this section, we introduce our categorization at the second level and the algorithms incorporated into BlackboxBench for each category at the third level. Regarding the method selection, we include classical techniques as baselines and advanced methods to underscore recent advancements in this field.
A-1 Categorization of implemented decision-based attacks
The primary aim of decision-based attacks is to devise effective strategies that render adversarial perturbations invisible. Existing methodologies can be categorized into two groups:
  • Random search methods These methods involve random direction updates and heuristic parameter adjustments. They include diverse strategies to acquire update directions, such as sampling from normal distributions, using evolutionary algorithms, gradient-free iterative techniques, and discrete optimization. (e.g., Boundary attack[45], Evolutionary attack[46], GeoDA[47], SFA[48] and Rays[49]).
  • Gradient-estimation-based methods These methods rely more on the information from the model feedback to guide the search direction. They transform decision-based attacks into optimization problems and incorporate gradient estimation, step size calculation, and binary searching as a framework. Subsequent methods enhance gradient estimation within this framework, and others introduce geometric patterns for improved adversarial detection. (e.g., OPT[50], Sign-OPT[51], HSJA [53], QEBA[54], NonLinear-BA[55], PSBA[56], Triangle attack[52]) and CISA[2].
A-2 Categorization of implemented score-based attacks
The objective of score-based attacks is to progressively move the adversarial sample closer to the decision boundary, allowing minimal input perturbations to mislead the model. Existing methods could be classified into three categories:
  • Random search methods These methods use stochastic strategies (e.g., sampling) to iteratively modify adversarial perturbations (e.g., SimBA[30], Parsimonious attack[31], Square attack[32], PPBA[33] and BABIES[34]).
  • Gradient-estimation-based methods These methods utilize model feedback or prior knowledge to approximate gradients for adversarial example creation (e.g., NES[35], Bandits[36], ZO-signSGD[37], SignHunter[38], AdvFlow[39] and NP-attack[40].
  • Combination-based methods These methods aim to enhance attack performance by integrating extra information from surrogate models with query feedback returned by the target model, to achieve fewer queries and higher attack success rate (e.g., Subspace attack[43], PRGF[1], BASES[44], $\mathcal{CG}$-Attack[41], and MCG[42]).
As a result, we implemented 29 query-based black-box attack methods in total, covering two types of query feedback. Every method is briefly described in Supplementary G-B and G-A.
A-3 Categorization of implemented transfer-based attacks
Amongst the earliest transfer-based attacks was the Iterative Fast Gradient Sign Method (I-FGSM)[81], which optimizes Eq. (5) iteratively, serving as a stronger adversary for white-box models compared to the single-step attack, $i.e.$, Fast Gradient Sign Method (FGSM)[91]. However, iterative attacks tend to overfit the surrogate model, causing poor transferability across models[61]. So can we achieve better trade-off between attack ability and transferability? Let $\mathcal{T}\left(\cdot\right)$ be the transformation function, $\mathcal{M}$ be the fusion operator, and I be some auxiliary information (specified later). $f^{\prime\prime}(x;\theta^{\prime\prime})\in\mathcal{F}$ indicates one refined surrogate model belonging to the surrogate ensemble $\mathcal{F}$, parameterized by $\theta^{\prime\prime}$. According to the general formulation:
$\arg\max_{\boldsymbol{x}^{*}}\mathcal{M}_{f^{\prime\prime}\in\mathcal{F}}(\mathcal{L}\left(f^{\prime\prime}\left(\mathcal{T}\left(\boldsymbol{x}^{*}\right);\theta^{\prime\prime}\right),y;\textbf{I}\right)),$ (A.1)
[…]
  • Data perspective Methods from this perspective use input transformation $\mathcal{T}\left(\cdot\right)$ on the benign sample to alleviate the overfitting to the surrogate model (e.g., PGD[57], TI-FGSM[58], SI-FGSM[59], Admix[60], DI2-FGSM[61], SIA[3]).
  • Optimization perspective Methods from this perspective introduce gradient-based optimization algorithms to escape from poor local maxima or modify the gradients to facilitate transferability (e.g., MI-FGSM[62], NI-FGSM[59], PI-FGSM[63], VT[64], RAP[65], LinBP[66], SGM[67], PGN[68]).
  • Feature perspective Methods from this perspective distort intermediate layer features by designing a new adversarial loss function $\mathcal{L}\left(\cdot;\textbf{I}\right)$ with the aid of auxiliary information I, such as the adversarial direction guide in ILA[69], the aggregate gradient serving as feature importance in FIA [70], and the integrated attention serving as neuron attribution in NAA[71].
  • Model perspective One main category focuses on model tuning: by refining the base surrogate model $f^{\prime}$ to a better one $f^{\prime\prime}$ (e.g., DRA[74], IAA[75], SWA[4]), or generating a set $\mathcal{F}$ of diverse variants from the base surrogate model (e.g., RD[72], GhostNet[73], LGV[72], Bayesian attack[4]), this category minimizes the misalignment between surrogate and target, thereby improving transferability. Another main category focuses on the fusion strategy modeled by $\mathcal{M}$: it reconciles the different gradients of an ensemble of surrogate models to capture the intrinsic transfer information (e.g., Logit ensemble[62], Loss ensemble[62], Longitudinal ensemble[73], CWA[76], AdaEA[77]).
As shown in Fig. 3 of the main text, apart from I-FGSM, BlackboxBench incorporates 29 additional transfer-based attacks, categorized into four groups. More detailed descriptions of transfer-based attacks can be found in Supplementary G-C. In addition to aforementioned individual transfer attacks, BlackboxBench supports the creation of customized composite attack pipelines through the flexible combination of existing functional blocks, each representing the core of one attack. BlackboxBench officially designs 98 pairwise composite attacks as examples, which will be discussed in Sec. IV-B.
A-4 Categorization of implemented defenses
  • Training-time approaches The most prominent training-time defense is Adversarial Training (AT) [93, 94, 95, 96, 97, 98], which is a defense strategy designed to bolster the robustness of ML models against adversarial attacks by incorporating adversarial examples during training. While this method enhances robustness, it demands more computational resources and may compromise accuracy on clean data. AT is deployed against both query-based attacks and transfer-based attacks in BlackboxBench.
  • Inference-time approaches Compared to traditional AT, defense deployed at inference stage (e.g., RND [99] and AAA [100]) are more lightweight and computationally efficient. Furthermore, the flexibility of inference-time approaches also allows them to be combined with other defense strategies, such as adversarial training, to harness the strengths of both. RND and AAA are mainly against query-based attacks in BlackboxBench. More descriptions of AT, RND, and AAA can be found in Supplementary H.
TABLE III: The ten analytical tools provided in the analysis module, organized into four functional categories. Note that Attack-component effect visualization category includes five distinct plot tools. For each tool, we describe its analytical purpose and indicate the specific key questions to which it is applied in the analysis of black-box adversarial vulnerability. The tool abbreviations used here ($e.g.$, AEV, MAV-A, DBV-A, etc.) correspond to those listed in the “Used tool” column of Tab. I.
Functional category Analytical tool Functionality Application to key question in Tab. I
[…]
Model attention visualization (MAV) A. Full-Gradient Saliency Maps (FullGrad) To visualize model attention in the spatial domain by highlighting pixel-level discriminative regions, enabling revealing how adversarial examples shift the model’s focus in the spatial domain. Attack budget-Q1: How about the methods’ reliance on the pertu- rbation norm budget? Attack procedure-Q1: How does the attacker traverse from the be- nign example to itsperturbed counterpart?
B. Frequency Saliency Map (FSM) To visualize model attention in the frequency domain by highlighting frequency-level discriminative components, enabling revealing how adversarial examples shift the model’s focus in the frequency domain. Defense-Q3: Are adversarial images transferable between normally trained models and adversarially trained models?
Adversarial divergence measurement (ADM) Conditional Adversarial Distribution (CAD) modeling & Adversarial Divergence metric To measure the distributional divergence of perturbations between models, providing a quantitative view of adversarial transferability. Model architecture-Q2: What kind of misalignment between surrogate models and target models then undermines transferability?
Decision boundary visualization (DBV) A. Decision Surface To visualize the proximity of benign images to the decision boundary, thus helping analyze the individual vulnerability to adversarial attacks. Data-Q3: How does the dimensionality the adversary restricted to affect its performance?
B. dbViz To visualize the structure of decision boundaries around benign examples on the data manifold, thus helping understand the boundary complexity.1 1 1 1 1 1 1 1 1 $111$ Data-Q2: Whether there are samples harder to be attacked by all attack methods? Model architecture-Q1: Which model architecture is most resistant to malicious queries?
B Codebase
B-1 Analysis module
To facilitate our understanding of black-box adversarial vulnerability, we introduce ten analytical tools grouped into four categories: attack component effect visualization, model attention visualization, adversarial divergence measurement, and decision boundary visualization. In the following, we elaborate the types of analyses we aim to conduct, the corresponding tools employed, and the underlying principles behind these tools. Tab. III summarizes these tools at a high level, while technical details provided in Supplementary I.
Attack-component effect visualization tools An adversarial attack scenario comprises four fundamental components: the method, attack setting, target model, and dataset. Nevertheless, the influence of each of these components remains unclear. To unravel the effect of each component, we fix other components, expand the one’s scope, and summarize the results via five plot tools (line charts, scatter charts, bar charts, bubble charts and heatmaps) to explore the underlying patterns.
Adversarial divergence measurement tools BlackboxBench aims to qualitatively interpret the adversarial transferability and establish a quantitative metric for evaluating transferability between white-box surrogate models and the target model. A high transferability implies that the perturbation crafted by the surrogate resembles that crafted by the target. To gain this resemblance, we adopt the Conditional Adversarial Distribution (CAD) modeling technique[41], which models the distribution of perturbations conditioned on clean examples. Based on the CAD, we design a transferability evaluation metric—Adversarial Divergence—to quantify the difference between models from an adversarial view by calculating the normalized Kullback-Leibler divergence between two CADs.
Decision boundary visualization tools To visualize the proximity of benign images to the decision boundary, we employ Decision Surface[103]. We also utilize dbViz[82] to visualize the information of decision boundary around benign images, aiding in understanding the difficulty for attackers trying to misclassify a benign image into wrong categories.
B-A Defense module
While our primary focus when building the BlackboxBench lies in implementing black-box attack algorithms, we also offer defense options for users to employ attacks with defense strategies. By integrating this module into the evaluation process, we aim to create a holistic understanding of both the robustness of defensive measures and the potency of attack strategies. This module is designed to rigorously assess the effectiveness of black-box attack algorithms when targeting neural network models equipped with defensive methods. We provide both static defense and dynamic defense in BlackboxBench. As the involvement of AT defenses is accomplished through the utilization of adversarially trained target models, our defense module exclusively provides three dynamic defense methods: RND[99], AAA-Linear[100], and AAA-Sine[100]. Our defense module offers a standardized testing environment, ensuring that evaluations are both consistent and replicable.
B-B Evaluation module
The evaluation module includes three crucial metrics: average query number (AQN), median query number (MQN), and attack success rate (ASR). AQN signifies the average query number over all images made during successful query-based black-box attacks. Similarly, MQN captures the middle value among all query numbers of successfully crafted adversarial examples. ASR measures the percent of adversarial samples classified as a different class in untargeted attacks or the target class in targeted attacks. For AQN and MQN, a lower value indicates a better attack performance, while for ASR, a higher value indicates a better attack performance.
C Additional Evaluation and Analysis
C-A Result overview
@@
  • Which task is more challenging, targeted or untargeted? Targeted attacks do pose a greater challenge than untargeted ones. In query-based attacks, regarding attacks with different targets, we can see that targeted query-based attacks require more query costs than untargeted attacks while exhibiting a relatively lower ASR. Similarly, achieving transferability in targeted attacks proves more challenging than in untargeted attacks. Despite our efforts to conduct evaluations with additional iterations and a logit loss following [104], along with a higher norm budget $\epsilon$ to boost targeted attack transferability, existing literature still shows its inability to fool models into predicting a targeted class, even though they perform well in untargeted attacks. Nevertheless, methods from the model-tuning perspective dramatically narrow the gap between targeted and untargeted tasks, illuminating a promising research direction to overcome the low transferability in targeted attack scenarios.
C-B Effect of Data
C-B1 Fluctuation among samples
Figure 21: Generalizability of evaluation results. Let random variable $X_{c}\in\mathcal{X}$ be the input image from class $c$ with $\mathcal{X}\subset\mathbb{R}^{d}$, and $x_{c}^{i}$ the realizations. Let random variable $A_{c}$ be the attack result on $X_{c}$, and $a_{c}^{i}$ the realizations. $A_{c}\subset\mathbb{Z}_{0}^{+}$ takes values in $[0,1]$ and $[0,M]$ when the attack result represents ASR and AQN, respectively, where $M$ is the maximum query number. This figure shows the averaged mean $\pm$ std of sample-level attack results over 5 random images sampled within one class. Here, averaged mean is defined as $\bar{\mu}=\sum_{c=1}^{C}\hat{\mu}_{c}/C$, averaged std is defined as $\bar{\sigma}=\sum_{c=1}^{C}\hat{\sigma}_{c}/C$, where $\hat{\mu}_{c}=\sum_{i=1}^{n}a_{c}^{i}/n$, $\hat{\sigma}_{c}=\sqrt{\sum_{i=1}^{n}\left(a_{c}^{i}-\hat{\mu}_{c}\right)^{2}/(n-1)}$ represent the mean and std of sample-level attack results within class $c$, respectively. In our trials, $n=5,C=1000$.
We believe that the observed stability in attack results considered on the entire dataset does not necessarily imply a low variation in attack results of single samples within a class [^7] The variance of dataset-level attack result is $Var(\sum_{c=1}^{C}A_{c}/C)=\sum_{c=1}^{C}Var(A_{c})/C^{2}$. Thus, the variance of dataset-level attack result is approximately $C$ times smaller than that of sample-level. Given that $A_{c}\in[0,N]$ (e.g., ASR of $X_{c}$$\in\{0,1\}$), the maximum of $Var(A_{c})$ is $N^{2}/4$. Thus the maximum variance of dataset-level attack result is $\sqrt{N^{2}/4C}$.. In contrast, Fig. 21, which plots the mean $\pm$ std of sample-level attack results averaged across classes, indicates a large fluctuation in attack results during the switch of samples.